(a) How much charge can be placed on a capacitor with air between the plates before it breaks down if the area of each plate is 4.00 cm2?

I am getting 10.26 micro Farads -- and being told this is incorrect. I am doing: K*E_0*A*Emax = 10.26 micro Farads

(b) Find the maximum charge if porcelain is used between the plates instead of air. nC

Have not attempted yet.
Answer #1

a)  The dielectric strength of air is not a fixed number. An accepted value is 3*10^6 V/m, The field in the capacitor is σ/ε where σ = Q/A. For air ε = ε0 = 8.85*10^-12 F/m The maximum charge is then

Q= 3*10^6*0.0004*8.85*10^-12 = 10.62*10^-9 C = 10.62 nC

(b) The dielectric strength of porcelain is 12*10^6; the dielectric constant is 6, so ε = 53.1*10^-12 F/m

Q = 12*10^6*0.0004*53.1*10^-12 = 254.88 nC